METU Department of Mathematics Math 304 History of Math Concepts II Spring 2011

Entrance Test February 17 2011

Name: Question 1. Which method of reasoning is called ”deduction” ? Which method of reasoning is called ”induction” ? Question 2. Which method is called ”reductio ad absurdum” (proof by contradiction) ? Question 3. Which process is called ”analysis” and which one ”synthesis” ? Question 4. What is your idea about ”Mathematical Analysis” as a branch of Mathematics ? Question 5. What is the difference between ”Theorems” and ”Lemmas” as mathematical claims ? Question 6. What does mean ”Corollary” ? Question 7. How to write 1696 in Roman numbers ?

Question 8. What does mean ”axioms” and why they are used in mathematics ? What does mean ”Axiomatic method” ?

Question 9. When (approximately) and where did live Euler, Pythagoras, Riemann, Fermat, Newton, and Euclid ?

Question 10. What is called ”hexagon” ? What the degree of a ”quartic equation” ? What is called ”octahedron” ?

Question 11. What do you know about ”Bourbaki” ?

Question 12. What do you know about mathematical Awards such as ”Fields medals” and the ”Millennium prises” ?

Question 13. How much is ”10110” in binary numeral system ?